Next-door neighbors Bob and Jim use hoses from both houses to fill Bob's swimming pool. They know that it takes 16 h using both hoses. They also know that Bob's hose, used alone, takes 50% less time than Jim's hose alone. How much time is required to fill the pool by each hose alone?
Answer by edjones(8007)

together 16 hr
Let jim's=j and bob's=.5j
1/j + 1/.5j = 1/16
8j+16j=.5j^2
24j=.5j^2
.5j^2-24j=0
j^2-48j=0 multiply each side by 2.
j(j-48)=0
j=48 hr
b=24 hr
.
Ed